OpenStudy (anonymous):

Help me find the quotient and simplify? m^2-n^2 m --------- ÷ -------- m-n m^2-nm So far I have it set up as "keep, change, flip". I can do this with numbers, it throws me off when there's only letters D:

OpenStudy (anonymous):

\[\frac{m^2-n^2}{m-n}\times \frac{m^2-nm}{m}\] is a start

OpenStudy (anonymous):

then factor a bunch then cancel

OpenStudy (anonymous):

"keep, change, flip"??!!

OpenStudy (anonymous):

I have that. I combine them and I have (m^2-n^2)(m^2-nm) ------------------ (m-n) for the top part would it be m^4-n^m ? And yes, keep change flip. You keep the first section, change multiplication to division and flip the last section.

OpenStudy (anonymous):

I'm just not sure if I'm combining correctly

OpenStudy (usarmy3947):

letsf take it piece by piece: m^2 - n^2 = (m - n)(m + n) m^2 + mn + n^2 = (m + n)(m + n) 7m^2 = 7*m*m nm^2 - mn^2 = nm(m - n) now lets put it all together: m^2 - n^2.............7m^2 --------------------- * -------------------- = m^2+2mn +n^2....nm^2 - mn^2 (m-n)(m+n)............7m*m ---------------------* ---------------- = (m+n)(m+n)........nm(m-n) (m-n)*(m+n)*7*m*m ---------------------------------- = (m+n)*(m+n)*n*m*(m-n) lets divide out the (m-n) (m+n)*7*m*m ------------------------- = (m+n)*(m+n)*n*m and then divide out the (m+n) 7*m*m ---------------- = (m+n)*n*m and last divide out m 7*m ------------ = answer (m+n)*n

OpenStudy (anonymous):

Awesome, thanks.

OpenStudy (usarmy3947):

no problem

OpenStudy (usarmy3947):

oh sorry Validavi Welcome to Open Study